# Benchmarking Mental State Representations in Language Models

### v1

@@ 3.3 Models
Zhu et al. (2024) have used two models for their experiments: Mistral-7B-Instruct (Jiang et al., 2023) and DeepSeek-7B-Chat (Bi et al., 2024) – both being the same size and fine-tuned. In contrast, we study two families of LMs that offer us options in model sizes and fine-tuning: Pythia (Biderman et al., 2023) and Llama-2 (Touvron et al., 2023). While Llama-2 offers “chat” versions fine-tuned using supervised learning and RLHF, Pythia’s open-source training set (Gao et al., 2020) ensures that there is no data leakage [^1] Llama-2 was released later than BigToM.. Additionally, we consider a version of Pythia-6.9B fine-tuned on a mixture of open-source instruction datasets (Wang et al., 2024), which we refer to as Pythia-6.9B-chat. [^2] https://huggingface.co/allenai/open-instruct-pythia-6.9b-tulu A summary of the models we study is reported in Table 2.
@@ 3.4 Probing Experiments
We aim to contribute to understanding how LMs represent beliefs of self and others by proposing a set of extensive probing experiments across LMs that differ in architecture, size, and fine-tuning approach. Our approach is generally similar to the one used by (Zhu et al., 2024), but we make a different operational choice: While (Zhu et al., 2024) train probes on each attention head for every layer, we train probes on the residual stream for every layer. We opted to use the residual stream as it integrates information from both the attention and feed-forward components, potentially encoding richer representations. Additionally, since the residual activations directly contribute to the final output predictions, probing them may better align with understanding the model’s behaviour for downstream tasks.
Model Size and Fine-tuning
We first report experiments to better understand the effect of model size and fine-tuning on belief probing accuracy. Specifically, we ask the following questions: Is there a relation between model size and probing accuracy? (RQ1) Does fine-tuning an LM with instruction-tuning or RLHF have an effect on probing accuracy? (RQ2) To answer these questions we performed the same probing experiment across all our models and compared the results.
Figure 2: Belief probing accuracy across models with different architecture, size and fine-tuning.
@@
Figure 4: To investigate potential memorisation in the probes, we compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components. For Llama2: All(7b) = $4096$, All(13b) = $5120$, All(70b) = $8192$. For Pythia: All(70m) = $512$, All(410m) = $1024$, All(1b) = $2048$, All(6.9b) = $4096$, All(12b) = $5120$. We report results for Protagonist beliefs. Results for Oracle are shown in Figure 8.
Table 1: Comparison of the effects of ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023) activation editing methods on three tasks from BigToM (Gandhi et al., 2023). TB denotes a true belief task, whereas FB denotes a false belief task. The numbers represent accuracy scores, with the difference in performance compared to no intervention (No int.) indicated as subscripts (ITI -- No int. and CAA -- No int.). An asterisk (*) denotes a statistically significant difference from No int. based on a McNemar’s test (McNemar, 1947) with $p<0.05$.
@@
TB FB TB $\land$ FB TB FB TB $\land$ FB TB FB TB $\land$ FB
@@
ITI $44_{+0}$ $44_{+0}$ $44_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
CAA $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $74_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+30}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
@@
ITI $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $69_{+0}$ $55_{+0}$ $37_{+0}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$
CAA $70_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}$ $72_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $69_{+0}$ $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $53_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $84_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}^{*}$

### v3

@@ 3.3 Models
We study two families of LMs that offer us options in model sizes and fine-tuning: Pythia (Biderman et al., 2023) and Llama-2 (Touvron et al., 2023) – for a total of 12 models. While Llama-2 offers “chat” versions first trained with SFT and then RLHF, Pythia’s open-source training set (Gao et al., 2020) ensures that there is no data leakage [^1] Llama-2 was released later than BigToM.. Additionally, we consider a SFT version of Pythia-6.9B trained on open-source instruction datasets (Wang et al., 2024), which we refer to as Pythia-6.9B-chat. [^2] https://huggingface.co/allenai/open-instruct-pythia-6.9b-tulu We provide model details in Table 2.
@@ 3.4 Probing Experiments
To study how LMs represent beliefs of self and others, we propose a set of extensive probing experiments across LMs that differ in architecture, size, and fine-tuning regime. We train probes on the residual stream, as it integrates information from both the attention and feed-forward components, potentially encoding richer representations. Additionally, since the residual activations directly contribute to the final output predictions, probing them may better align with understanding the model’s behaviour for downstream tasks.
Control Tasks
Depending on the model dimension, the probes we train have a significant number of learnable parameters – up to $16,385$ for Llama-2-70B. This raises the concern that probes might learn to rely on irrelevant patterns in the data instead of capturing meaningful relationships. To account for the potential confounding effect of hidden state size, we include two controls. First, following Hewitt and Liang (2019), we train and evaluate probes on a version of $\mathcal{D}^{P}_{p}$ with randomly permuted labels – thus removing real input-label relationships. If a probe still performs well on the permuted data, this suggests it may be exploiting superficial correlations rather than capturing genuine structure. Second, we effectively reduce the number of learnable parameters in the probes by projecting $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ onto their $k$ largest principal components using PCA before training. This minimises the risk of the probes relying on spurious patterns in the data.
Robustness Tests
Previous work left the impact of prompting on belief probing accuracy unexplored. Our second set of experiments aims to study whether belief representations are robust to different prompts. Research on prompt robustness in language models focused mainly on revealing vulnerability to prompt alterations on downstream performance (Min et al., 2022; Ishibashi et al., 2023; Shaikh et al., 2023; Leidinger et al., 2023; Sclar et al., 2024). In contrast, we study how different prompt alterations influence probing performance, i.e. models’ internal representations. Unlike model outputs that are shaped by decoding strategies, which act as confounders, models’ activations are more abstract and offer a better lens into how robust or brittle internal representations are. We define four prompt variations:
  • Random: Following Gurnee and Tegmark (2024), we add 10 random tokens to the belief statement.
  • Time Specification: The prompt specifies that the belief statement refers to the end of the story. We include this variation because some belief statements can be true (false) at the story’s beginning but false (true) at the end. For example, consider the story in Figure 2: if Noor does not witness the swap, in the end, she will believe the pitcher contains almond milk ($z_{p}=\mathrm{True}$). However, if the same belief is referred to the beginning of the story, then it is false ($z_{p}=\mathrm{False}$).
While all maintain conceptual and semantic parity with the Original prompt used in (Zhu et al., 2024), Random and Misleading are expected to negatively impact LMs’ representations, while Time Specification and Initial Belief are supposed to have a positive influence. Robust representations of beliefs should exhibit minimal sensitivity to these alterations. Our experiments compare probe accuracy across different model sizes, fine-tuning, and prompt variations. Examples of prompts are reported in Section A.1.4.
Figure 3: Belief probing accuracy show similar patterns across all models: oracle belief representations generally form already in the first layers, while protagonist belief representations emerge at the intermediate layers. Moreover, probing accuracy increases with model size and, more crucially for smaller models, with fine-tuning.
3.5 Activation Editing
Prior work found that it is possible to manipulate models’ representations of beliefs by using (Li et al., 2023b, ITI), and that such interventions can improve LMs’ performance on ToM tasks. We take this further by asking whether a general “belief vector” can be distilled and injected into the models’ activations to strengthen their ToM abilities. To this end, we use contrastive activation addition (Rimsky et al., 2023, CAA), an extension of activation addition (Turner et al., 2023, AA) that computes steering vectors to control LMs’ behaviour. Steering vectors are computed as the average difference in residual stream activations between pairs of positive and negative instances of a specific behaviour. Formally, given a dataset $\mathcal{D}$ of triplets ($p$, $c_{p}$, $c_{n}$), where $p$ is a prompt, $c_{p}$ is a positive completion, and $c_{n}$ is a negative completion, CAA computes a mean difference vector $v^{md}_{l}$ for layer $l$ as:
$v^{md}_{l}=\frac{1}{\absolutevalue{\mathcal{D}}}\sum_{p,c_{p},c_{n}\in\mathcal{D}}a_{l}(p,c_{p})-a_{l}(p,c_{n})$ (1)
For example, in Figure 2, $p$ is the Story, $c_{p}$ could be the true belief, and $c_{n}$ the false belief. During inference, these steering vectors are multiplied by an appropriate coefficient $\alpha$ and added at every token position of the generated text after the prompt. CAA has two main advantages over ITI: First, it eliminates the need to train probes, making it computationally cheap. For example, for Llama2 70B, ITI needs to train 5,120 probes while CAA only needs to compute 80 vectors. Second, it operates at the residual stream level, making it easier to use than methods that intervene on specific attention heads like ITI. While CAA has been used to control alignment-relevant behaviour, such as hallucinations, refusal, and sycophancy (Rimsky et al., 2023), we are the first to apply it to enhance LMs’ ToM reasoning. The “belief vectors” (i.e. steering vectors) we obtain can be understood as isolating the direction in the LMs’ latent space corresponding to taking the perspective of another agent. To evaluate both base and fine-tuned LMs, we rank their answers to the ToM questions according to $p_{LM}(a|q)$ (Petroni et al., 2019). For a fair comparison, we adopt the train/test Forward Belief split used in (Zhu et al., 2024) to compute and evaluate the steering vectors. Additionally, we evaluate the transferability of the CAA steering vectors by applying them to two other BigToM tasks: Forward Action and Backward Belief. We provide details about these tasks in Section A.1.1, and a more detailed explanation ITI in Section A.5.
4 Results
@@
Figure 4: We compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $k=\{2,10,100,1000\}$ principal components. Results are for protagonist beliefs (for oracle see Figure 12). In general, it is possible to recover most of the original accuracy by training probes on a smaller number $k$ of principal components of the activations.
Figure 5: Sensitivity of protagonist belief probing accuracy to different prompt variations. Results for Pythia are shown in Figure 9. Representations are brittle to prompt variations.
Effect of Model Size and Fine-tuning
Results from our study on model size and fine-tuning are shown in Figure 3. For oracle beliefs, probing accuracy rapidly converges to $100$, with larger models showing faster convergence. Even the smallest Pythia-70m achieves $95\%$ accuracy. For protagonist beliefs, we notice a similar pattern across most models, where accuracy at early layers is particularly low and then increases at the intermediate layers. What happens at early layers is overfitting, which may be caused by spurious features introduced by the initial coding strategy of language models, where individual token representations are mixed together (Gurnee et al., 2023). We further discuss this in Section A.2.1.
In general, probing accuracy increases with model size, although there is a performance gap between Llama-2 and Pythia. For example, Llama2-13B reaches around $80\%$ accuracy, while Pythia-12B achieves approximately $60\%$. This gap is likely due to Llama-2 being trained on nearly seven times more tokens than Pythia (cf. Table 2). Probes from fine-tuned LMs show significantly better accuracy, with improvements of up to $+29\%$ for Llama2-7B-chat (SFT + RLHF) and $+26\%$ for Pythia-6.9B-chat (SFT) compared to probes from their base version. The same probes outperform (Llama-2) or are on par (Pythia) with probes trained on twice as large base models (12/13B). This highlights a key role of fine-tuning in shaping belief representations in smaller LMs. The performance gap closes for the largest Llama2-70B, for which the improvements from fine-tuning are marginal.
We characterise the relationship between probe accuracy and model size in Figure 7, using the best accuracy for each LM – i.e., the highest accuracy among probes ${g_{l}}$ trained on activations ${a_{l}}$ for model $f$. For Llama-2 base and Pythia base, probing accuracy scales logarithmically with model size (Figure 7(b), 7(d)), while for fine-tuned Llama-2 models, it scales linearly (Figure 7(c)).
Control Tasks
Figure 8(a) shows that probes trained on the control task consistently perform at random chance, confirming that higher probing accuracy in larger models meaningfully reflects a greater ability to extract ToM representations, rather than simply being a by-product of spurious correlations. For Llama models, the probes generally exhibit selectivity: they achieve high accuracy when probing for beliefs but remain at chance level on control tasks. Pythia’s overall accuracy is too low to allow for selectivity.
Figure 4 shows probing accuracy on protagonist when training the probes on the top $k$ principal components of Llama-2’s internal activations. We provide results for Pythia in Figure 11, and for all models on oracle settings in Figure 12. We consider $k=\{2,10,100,1000\}$, spanning several orders of magnitude. [^3] For models with hidden dimensions smaller than $1000$, we skip this value. Results show that it is generally possible to recover most of the original accuracy by training probes on a smaller number $k$ of principal components of the activations. We also performed the first control experiment, this time only using the first $k=\{100,1000\}$ principal components. Figure 8(b) and 8(c) again show that probes trained on the control task consistently perform at random chance, confirming that probes are not fitting spurious patterns. Additionally, this suggests that belief representations are embedded in a low-dimensional subspace $\mathcal{B}$ spanned by the top $k$ eigenvectors $\{v_{1},\dots,v_{k}\}$ of the covariance matrix $\mathsf{C}=\mathbb{E}[(a-\mathbb{E}[a])(a-\mathbb{E}[a])^{\top}]$.
Sensitivity to Prompting
Figure 5 compares protagonist probe accuracy across various prompt variations for Llama-2 models. As can be seen from the figure, providing the protagonist’s Initial Belief in the story yields higher probe accuracy compared to the Original prompt. Accuracy for all the other prompt variations is generally lower than Original. Misleading prompts hurt performance across all models. This finding resonates with Webson and Pavlick (2022), who found that instruction-tuned models, despite being more robust, are still sensitive to misleading prompts. On the other hand, Time Specification unexpectedly does not help in disambiguating belief states in different time frames, as we hypothesised in §3.4. Additionally, models show sensitivity to Random tokens placed before the belief statement. Pythia models show similar patterns, shown in Figure 9. Results for oracle beliefs are reported in Figure 10 and indicate that models maintain high accuracy. Misleading prompts slightly reduce performance to around $95\%$. In summary, these experiments show that LMs possess robust belief representations when taking an omniscient perspective, whereas their representations of others’ beliefs are brittle to prompt variations.
@@
TB FB Both TB FB Both TB FB Both
@@
ITI $44_{+0}$ $44_{+0}$ $44_{+0}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$
CAA $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}^{*}$ $71_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+27}}^{*}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}^{*}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}^{*}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}^{*}$ $74_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+30}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$
@@
ITI $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $69_{+0}$ $55_{+0}$ $37_{+0}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+3}}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$
CAA $70_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+14}}$ $72_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}^{*}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $69_{+0}$ $67_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $53_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}$ $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $84_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+27}}^{*}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}^{*}$